# Heralded Emission Detection in Quantum Dot Solids under Twin-Photon Excitation

### v1

@@ Abstract
Harnessing quantum correlations between photons is an emerging direction in optical spectroscopy, but experimental progress has thus far been limited to fluorescence-detected studies of molecular systems at room temperature. Here, we demonstrate heralded emission detection (HED) under continuous-wave (CW) entangled-photon excitation of near-infrared (NIR)–emitting colloidal III–V quantum dots (QDs) at low temperatures. We use superconducting nanowire single-photon detectors (SNSPDs) and large-area NIR avalanche photodiodes (APDs) to maximize the time resolution ($\sim$70 ps) and emission count rates ($\sim$2000 cps). To our knowledge, these results represent the first demonstration of HED in the solid state and establish NIR QDs as a bright, tunable model system with longer-lived coherences and integrability with optical cavities as an additional experimental control knob.
1 Introduction
@@
Optical spectroscopy continues to illuminate foundational processes in chemical, biomolecular, and solid-state systems across timescales. Thus far, canonical optical spectroscopy uses classical light from pulsed or continuous-wave (CW) lasers producing coherent wavepackets or wavetrains. The conventional semiclassical description for the light-matter interaction, treating light as a classical perturbation of a quantum system, is therefore typically sufficient [1, 2]. Electronic or vibrational coherences generated through pulse/sample interaction can be interfered with additional time-delayed optical pulses, thereby bestowing non-linear spectroscopy with exquisite control via the bandwidth, shape, and delay between pulses [3]. These control knobs can extract the energy-time correlations of a sample response without requiring any quantization of field excitations, i.e., without invoking photons.
@@
A relatively new paradigm in optical spectroscopy uses quantum light i.e., photons with defined correlations in time, energy, or polarization [4, 5]. This is distinctly different from classical spectroscopy, where coherent laser light is characterized by a Poissonian number distribution of otherwise uncorrelated photons [6, 7]. Quantum spectroscopy can therefore use what classical spectroscopy cannot; the inherent correlations between photons in bespoke quantum states. Recent theoretical efforts have explored non-classical correlations between photons in spectroscopic experiments e.g., using energy-time entangled pairs. Intriguing theoretical predictions [4] beyond overcoming classical noise limits have been made [8]. For instance, access to classically forbidden excitation pathways [9], control over photo-chemical conversion rates through vibrational wavepacket preparation [10, 11], or remote-controlled spectroscopy via twin-pair detection of entangled states have been proposed [12].
@@
Only a few experimental studies on entangled-photon light-matter interaction follow the theoretical lead. As such, the optical response of materials and molecules under quantum light excitation, e.g., single-photon Fock states or energy-time entangled photon pairs from spontaneous parametric downconversion (SPDC) has gradually come into focus [13, 14], but is still limited by the significant experimental complexity of quantum light spectroscopy. In this context, the still-limited brightness of entangled photon sources used for sample excitation is a particular challenge [15]. This is different from most spectroscopic experiments, in which sample stability rather than the achievable excitation power is the limiting factor. Probing sample coherences after entangled-photon interaction with the sample using a ’probe’ pulse (or second photon) is therefore not easily achieved due to the low probability of two photons interacting with the same emitter. Instead, either the entangled photon phase delay (in transmission) or spontaneous emission as proxy for entangled photon absorption is measured.
@@
Two common entangled two-photon (ETP) spectroscopies include fluorescence-detected two-photon absorption [16, 17, 18] and heralded emission detection (HED) [19, 20, 21, 22, 23]. Two-photon absorption using ETP has been reported to exhibit stronger absorption compared to classical two-photon absorption, owing to the intrinsic time correlation of photons in pairs [16]. This effect has potential applications in fluorescence microscopy [17], although it remains somewhat debated [24]. HED leverages the intrinsic time correlation of photon pairs from quantum light sources such as SPDC. One photon excites the emissive sample, while its twin is directly detected to “herald” the excitation event. Correlating the herald with the following photoluminescence photon reconstructs the emission dynamics, using the twin-pair correlations. This is fundamentally different from the laser-pulse correlations employed in conventional lifetime measurements using time-correlated single-photon counting (TCSPC). HED lifetime measurements therefore work with pulsed and CW lasers. The wavelengths of the entangled photons can be finely tuned by adjusting the temperature of the nonlinear crystal in the SPDC process, allowing for wavelength-dependent excitation studies. Because SPDC produces photon statistics in each arm that follow a thermal distribution, HED can be employed to study emission lifetimes under thermal light-like conditions. This makes it especially valuable for exploring systems such as photosynthetic complexes or photovoltaic materials under illumination with realistic underlying photon statistics. Another key difference is that TCSPC requires very low detection probability per pulse (1–5%) to avoid pile-up, which limits usable counts [25]. HED, not tied to an excitation clock, avoids pile-up and can in principle operate at higher fluxes. Despite its conceptual simplicity, only a few studies have applied HED, all of which have measured solution-phase molecular systems such as Indocyanine Green (ICG) [19], Rhodamine 6G (R6G) [20], and IR-140 [23]. A noteworthy feature of HED is its ability to herald single-photon Fock states to study light-driven processes in the ultimate limit of excitation intensity, e.g., in Light-Harvesting 2 (LH2) complexes [21, 22].
@@
To advance quantum spectroscopy, both the scope of accessible methods and samples must be expanded. HED’s current main limitation, shared with all other entangled-pair spectroscopy, is the low photon flux and coincidence rates from entangled-photon sources. HED is therefore currently not competitive for routine lifetime measurements. However, it can provide a foundation for extending quantum-light spectroscopy beyond simple intensity correlations, opening the possibility to resolve additional energy- and polarization-correlations. Such capabilities would enable more sophisticated approaches for probing quantum systems, with the broader aim of accessing entanglement signatures and unraveling excitonic dynamics in quantum materials [26]. In light of these requirements, the expansion of entangled-pair spectroscopy from solution-phase organic molecules to solid-state materials is a logical next step. First, the entangled light-matter interaction and emission can be photonically enhanced to increase both the overall brightness and excitation rate per emitter. The latter is needed, for example, to validate theoretical predictions of inter-pair delay-dependent two-photon excitation i.e., electronic pathway selectivity [27]. To advance quantum spectroscopy, both the scope of accessible methods and samples must be expanded. HED’s current main limitation, shared with all other entangled-pair spectroscopy, is the low photon flux and coincidence rates from entangled-photon sources. HED is therefore currently not competitive for routine lifetime measurements. However, it can provide a foundation for extending quantum-light spectroscopy beyond simple intensity correlations, opening the possibility to resolve additional energy- and polarization-correlations. Such capabilities would enable more sophisticated approaches for probing quantum systems, with the broader aim of accessing entanglement signatures and unraveling excitonic dynamics in quantum materials [26]. In light of these requirements, the expansion of entangled-pair spectroscopy from solution-phase organic molecules to solid-state materials is a logical next step. First, the entangled light-matter interaction and emission can be photonically enhanced to increase both the overall brightness and excitation rate per emitter. The latter is needed, for example, to validate theoretical predictions of inter-pair delay-dependent two-photon excitation i.e., electronic pathway selectivity [27].
@@
Here, we demonstrate HED in solid-state materials at low temperatures, specifically thin-films of near-infrared (NIR) emitting colloidal InAs/ZnSe quantum dots (QDs). QDs offer a rich platform for exploring entangled-pair light–matter interactions, thanks to their tunable electronic fine structure, phonon coupling, and electron–hole radiative and non-radiative rates [28]. To our knowledge, our results represent the first demonstration of entangled-photon heralded emission detection in a solid-state material, distinct from prior two-photon excitation approaches using entangled photons [29]. QD’s processability and proven integration with nanophotonic structures may help overcome brightness limitations through enhancement of the absorption and emission [30], and serve as a model system for entangled-photon non-linear spectroscopy.
@@ 2 Result
We developed an improved HED apparatus capable of operating over variable solid-state sample temperatures, as shown in Fig. 1a. A detailed description of the setup components is provided in the Supporting Information. Briefly, entangled photon pairs are generated via SPDC by pumping a periodically poled potassium titanyl phosphate (ppKTP) crystal with a 405 nm CW laser under the type-II phase-matching condition. The generated pairs exhibit a bandwidth of $\sim$2 nm with spectral overlap consistent with degenerate SPDC (Fig. 1b). A Hong–Ou–Mandel (HOM) visibility of 86% confirms the high indistinguishability of the photon pairs, in agreement with previous reports (Fig. S1) [31]. The two photons are separated by a fiber-coupled polarization beam splitter (PBS), yielding a heralding arm and a sample arm. The heralding photons are directly detected by an SNSPD, while the sample-arm photons are sent through a 50:50 beamsplitter into a cryostat with a 0.9 numerical aperature (NA) cryo-objective to excite the sample. The resulting emission is collected by the same objective, filtered from the excitation photons, and coupled via a fiber collimator into a second SNSPD. The use of SNSPDs is a notable feature of our apparatus due to their high photo-detection efficiency in the NIR ($\sim$80%), minimal dark counts (<<10 Hz), and low dead time ($\leq$35 ns). Their additionally low timing jitter (31 ps and 50 ps) provides an overall narrow ($\sim$72 ps) instrument response function (IRF) for coincidence counts without sample (Fig. 1c). The IRF width is mainly determined by the convolution of the timing jitter of the SPDC source, SNSPDs, and the single-photon counting module (19 ps).
We selected colloidal InAs/ZnSe core–shell QDs as the model system for HED. These QDs are promising NIR emitters[32] and are also widely applied in photonic devices such as photoconductors[33], phototransistors[34], and photodiodes[35, 36]. Importantly, colloidal InAs QDs can be directly deposited onto photonic platforms[37], offering a straightforward route toward integration with quantum photonic structures for optical communication and emerging quantum photonic applications.
@@
Our synthesized type-I heterostructured InAs/ZnSe core–shell nanocrystals display high (approximately 58%) ensemble photoluminescence quantum yield (PLQY) in solution, consistent with previously reported values for this material system [38]. This PLQY is substantially higher than of organic dyes in the NIR [39], rendering NIR-QDs a natural choice for HED with high signal strengths. Room temperature QDs ensemble absorption and emission spectra reveal a low prominence excitonic shoulder around 899 nm based on Elliott model fitting (Fig. S3) and a redshifted, broad emission around 984 nm, shown in Fig. 1d. We attribute the absence of narrower features to inhomogeneous broadening due to size- and shape-inhomogeneity, which is common even for the currently best InAs QDs [32, 40]. The transmission electron microscopy (TEM) image, as shown in the inset of Fig. 1d, confirms the tetrahedral QD shape [38] and a mean particle size of 8.2 $\pm$ 2.2 nm (Fig. S3). This size is only a fraction of the InAs bulk Bohr radius of around 35 nm [41], rendering our QDs well within the strong confinement regime, where size dispersion has a high impact on spectral broadening[32].
For emissive thin-film preparation, we employed drop-casting onto gold-coated coverslips (Fig. 1a) rather than conventional spin coating. Spin-coated films typically yield thin layers ($\sim$30–100 nm) [42, 43], which contain only a limited number of QDs within the optical focal volume (Rayleigh length × focus spot size), thereby reducing the probability of photon absorption. Consequently, the detected HED signals can be weak even for materials with high PLQY. One strategy to increase film thickness is to mix QDs with polymers such as polystyrene or poly(methyl methacrylate). However, achieving a thickness comparable to the Rayleigh length requires a high polymer concentration, which increases the polymer-to-QD ratio within the focal volume, reduces film brightness, and may ntroduce non-negligible quenching effects (see Supporting Information). In principle, higher QD loadings could offset this issue, but in practice such concentrations are difficult to achieve. Another potential approach is layer-by-layer deposition, in which oleic acid–capped QDs are sequentially coated, the surface wetting properties are modified via solid-state ligand exchange, and additional layers are deposited. While this method can produce multilayered QD films [44], it often comes at the expense of reduced PLQY and requires further optimization [43]. By contrast, drop-casting naturally produces thicker and denser films, thereby improving absorption efficiency and yielding stronger HED signals. The drawback, however, is that drop-cast films typically exhibit non-uniform coverage, as shown in Fig. S3. Representative SEM images, also in Fig. S3, further highlight their inhomogeneous morphology. This inhomogeneity makes it challenging to systematically study the relationship between film brightness and thickness.
@@
Figure 2: Data analyses in heralded emission detection. (a) In heralded emission detection (HED), photon pairs are split into a heralding arm (blue) and a sample arm (orange). Pairs can be classified as coincidence pairs, which are time-correlated and produce bunching in the $g^{(2)}(\tau)$, or non-coincidence pairs, which contribute uncorrelated background. The sample emission therefore possesses a similar statistics with emission photon from both twins (red) and accidental pairs (light red). (b) Intensity traces of the heralding arm and the emission signal during a 20h measurement at 4.7 K, demonstrating stable down-conversion over a ten-hour period. (c) Normalized cross-correlation $g^{(2)}(\tau)$ between the heralding arm and the emission at 4.7 K. The orange-shaded region corresponds to contributions from correlated heralding–emitted photon pairs (entangled two-photon down-conversion), while the purple region represents uncorrelated background with $g^{(2)}(\tau)=1$. (d) Schematic for the $g^{(2)}$ correlation analysis of time-tagged HED data. Delaying one sequence relative to the other only shifts the bunching peak without changing its shape. (e) Schematic of the effect of adopting a ’start–stop’ alternative from time-correlated single-photon counting (TCSPC) to HED, introducing distortions.
@@
Fig. 2 illustrates the concept, stability of emission, and different approaches for analyzing HED data. Three photon “arms” are considered in the measurement: the heralding arm, the sample arm used to excite the sample, and the emission arm, which collects photons emitted by the sample following excitation. SPDC sources are not ideal [45]; in addition to generating true coincidence photon pairs, non-coincidence pairs are also produced, as shown in Fig. 2a. Consequently, the emitted sample photons originate from both non-coincidence and coincidence pairs. The former produces accidental heralding-emission coincidences. The cample emission count rates remained stable over the course of the experiment ($\sim$15–18 hours), confirming the sample’s stability under low excitation fluences (see Fig. 2b for 4.7 K and Fig. S4 for all temperatures). At 4.7 K, the heralding arm consistently recorded an average count rate of $\sim 3.2\times 10^{6}$ counts per second (cps). Assuming the sample arm exhibits a similar rate, and accounting for losses through optical elements, the number of excitation photons reaching the sample is estimated to be $1.12\times 10^{6}$, corresponding to a photon power of $\sim$275 fW. Emitted photons were collected at a rate of $\sim$89 cps (see Supporting Information for all temperatures and details).
@@
We now analyze correlations between heralding and sample photons, extending approaches previously applied to molecular dyes at room temperature [19, 21, 20, 23, 22]. Leveraging the short 70 ps IRF of our SNSPDs, we resolve the normalized cross-correlation $g^{(2)}(\tau)$ between emission and heralding photons (Fig. 2c). As expected, a distinct bunching peak >1>1 at $\tau=0$ indicates preservation of the bunched statistics of twin pairs in the heralded emission. The background with $g^{(2)}(\tau)=1$ results from accidental coincidence events. The decay of $g^{(2)}(\tau)>1$ with $\tau$ measures the effective delay between heralded photon absorption and emission, i.e. the spontaneous emission lifetime [19]. We confirm the near-equivalency of HED lifetimes of our QDs and TCSPC, as shown in the Supporting Information. Small deviations are most likely due to differences in the excitation wavelength or small deviations in setup alignments between different excitation sources. Interestingly, we identify a consistent dip of $g^{(2)}<1$ for $\tau<0$ (Fig. 2c, inset). Such anti-correlations would be consistent with time reversal symmetry breaking, but likely have a more trivial origin. We assign the origin of this dip to a non-obvious artifact from detector deadtime. Coincidences with $\tau<0$ are from emitted photons leading heralding photons. A leading emitted photon has a larger than random probability of being led by a second, even earlier heralding photon. The detection of this second, earliest, heralding photon turns the heralding arm SNSPD dark deadtime (<<20 ns). The number of following emission/heralding coincidences within the deadtime is thus reduced, producing the dip in $g^{(2)}(\tau<0)$. This artifact may be misinterpreted or may go unnoticed, especially for longer deadtime detectors. While the above correlation analysis is established, we like to point out that translation of the “start-stop” paradigm used in pulsed-laser TCSPC would require correction in post-processing and is more prone to distortion from pathlength differences. We clarify this in Fig. 2d-e.
@@
The start–stop algorithm selects a photon on one channel as “start” and a photon on the other channel as “stop,” storing their time differences in a histogram. By contrast, the correlation algorithm accumulates all time differences between detection events on the two channels into a histogram, treating every click as both “start” and “stop,” and calculating both positive and negative delays. Using a start–stop scheme in HED can introduce distortions, since the overall number of pairs along the lifetime not only depends on the heralding ratio and entanglement time, but also the waiting time distribution of the accidental pairs. This can produce spurious accidental coincidences that are difficult to separate from the lifetime decay. Timing misalignment between emitted photon and their heralding photon can further deform the decay dynamics (Fig. 2d). The $g^{(2)}$ correlation approach avoids these complications, as it does not impose fixed start/stop assignments on the photons. Any time delay between the heralding and emission arms merely shifts the correlation peak without distorting its decay dynamics (Fig. 2e). Waiting time distributions between accidental counts are naturally factored out in the normalized $g^{(2)}$.
@@
One shortcoming of many SNSPDs is their single-mode fiber access [46], causing substantial insertion losses. If high count rates rather than minimal timing jitter are desirable in HED, large area APDs are preferred. Fig. 3a–c showcases HED results acquired across different temperatures using both an SNSPD (Fig. 3a) and a large-area APD (Fig. 3b-c). Although the employed APD exhibited high dark counts ($\sim$1848 cps), they contribute only to accidental coincidences and do not affect the shape of the bunching peak. We consistently observed total emission count rates of $\sim$4000 cps when accounting for accidental dark counts. Sample 1, with a quantum dot size of 8.2 nm, and Sample 2, with a size of 6.3 nm, showed similar absolute brightness despite their differing lifetimes. These detected count rates led to an increased external quantum efficiency (EQE) for the film i.e., the ratio of detected emission photons to heralded sample arm photons (Fig. 3f).

### v2

# Heralded Emission Detection in InAs/ZnSe Quantum Dot Solids Using Time-Correlated Photons

@@ Abstract
Harnessing quantum correlations between photons is an emerging frontier in optical spectroscopy, yet experimental demonstrations have largely remained limited to molecular systems at room temperature. Here, we investigate heralded emission detection (HED) under continuous-wave entangled-photon excitation of near-infrared (NIR)–emitting colloidal III–V quantum dots (QDs) solids at low temperatures. We demonstrate the advantages of superconducting nanowire single-photon detectors (SNSPDs) for high time resolution ($\sim$72 ps) and large-area NIR avalanche photodiodes (APDs) for high emission count rates ($\sim$2000 cps). Second-order photon-correlation analysis reveals exciton lifetimes and fine-structure energy splittings. These results establish NIR colloidal QDs as a bright, tunable model system for quantum-light spectroscopy and highlight their compatibility with optical cavities as a further experimental control parameter.
@@
Introduction Optical spectroscopy continues to illuminate foundational processes in chemical, biomolecular, and solid-state systems across timescales. Thus far, canonical optical spectroscopy uses classical light from pulsed or continuous-wave (CW) lasers producing coherent wavepackets or wavetrains. The conventional semiclassical description for the light-matter interaction, treating light as a classical perturbation of a quantum system, is therefore typically sufficient [scully1997quantum, boyd2008nonlinear]. Electronic or vibrational coherences generated through pulse/sample interaction can be interfered with additional time-delayed optical pulses, thereby bestowing non-linear spectroscopy with exquisite control via the bandwidth, shape, and delay between pulses [weinacht2018time]. These control knobs can extract the energy-time correlations of a sample response without requiring any quantization of field excitations, i.e., without invoking photons.
@@
The use of quantum light, i.e. photons with defined correlations in time, energy, or polarization [dorfman2016nonlinear, mukamel2020roadmap] is an emerging new paradigm in optical spectroscopy. This is distinctly different from classical spectroscopy, where coherent laser light is characterized by a Poissonian number distribution of otherwise uncorrelated photons [fox2006quantum, gerry2023introductory]. Quantum spectroscopy can therefore use what classical spectroscopy cannot; the inherent correlations between photons in bespoke quantum states. Recent theoretical efforts have explored non-classical correlations between photons in spectroscopic experiments e.g., using energy-time entangled pairs. Intriguing theoretical predictions [dorfman2016nonlinear] beyond overcoming classical noise limits have been made [aasi2013enhanced]. For instance, access to classically forbidden excitation pathways [muthukrishnan2004inducing], control over photo-chemical conversion rates through vibrational wavepacket preparation [gu2021photoisomerization, zhou2025enhanced], or remote-controlled spectroscopy via twin-pair detection of entangled states have been proposed [fujihashi2023probing].
@@
Only a few experimental studies on entangled-photon light-matter interaction follow the theoretical lead. As such, the optical response of materials and molecules under quantum light excitation, e.g., single-photon Fock states or energy-time entangled photon pairs from spontaneous parametric downconversion (SPDC) has gradually come into focus [couteau2018spontaneous, anwar2021entangled], but is still limited by the significant experimental complexity of quantum light spectroscopy. In this context, the still-limited brightness of entangled photon sources used for sample excitation is a particular challenge [tsao2025enhancing]. This is different from most spectroscopic experiments, in which sample stability rather than the achievable excitation power is the limiting factor. Probing sample coherences after entangled-photon interaction with the sample using a ’probe’ pulse (or second photon) is therefore not easily achieved due to the low probability of two photons interacting with the same emitter. Instead, either the entangled photon phase delay (in transmission) or spontaneous emission as proxy for entangled photon absorption is measured.
@@
Two common entangled two-photon (ETP) spectroscopies include fluorescence-detected two-photon absorption [burdick2021enhancing, varnavski2020two, gu2020manipulating] and heralded emission detection (HED) [harper2023entangled, eshun2023fluorescence, li2023single, alvarez2025correlated, gabler2025benchmarking]. Two-photon absorption using ETP has been reported to exhibit stronger absorption compared to classical two-photon absorption, owing to the intrinsic time correlation of photons in pairs [burdick2021enhancing]. This effect has potential applications in fluorescence microscopy [varnavski2020two], although it remains somewhat debated [hickam2022single]. HED leverages the intrinsic time correlation of photon pairs from quantum light sources such as SPDC. One photon excites the emissive sample, while its twin is directly detected to “herald” the excitation event. Correlating the herald with the following photoluminescence photon reconstructs the emission dynamics, using the twin-pair correlations. This is fundamentally different from the laser-pulse correlations employed in conventional lifetime measurements using time-correlated single-photon counting (TCSPC). HED lifetime measurements therefore work with pulsed and CW lasers. The wavelengths of the entangled photons can be finely tuned by adjusting the temperature of the nonlinear crystal in the SPDC process, allowing for wavelength-dependent excitation studies. Because SPDC produces photon statistics in each arm that follow a thermal distribution, HED can be employed to study emission lifetimes under thermal light-like conditions. This makes it especially valuable for exploring systems such as photosynthetic complexes or photovoltaic materials under illumination with realistic underlying photon statistics. Another key difference is that TCSPC requires very low detection probability per pulse (1–5%) to avoid pile-up, which limits usable counts [Lakowicz2006]. HED, not tied to an excitation clock, avoids pile-up and can in principle operate at higher fluxes. Despite its conceptual simplicity, only a few studies have applied HED, all of which have measured solution-phase molecular systems such as Indocyanine Green (ICG) [harper2023entangled], Rhodamine 6G (R6G) [eshun2023fluorescence], and IR-140 [gabler2025benchmarking]. A noteworthy feature of HED is its ability to herald single-photon Fock states to study light-driven processes in the ultimate limit of excitation intensity, e.g., in Light-Harvesting 2 (LH2) complexes [li2023single, alvarez2025correlated].
@@
To advance quantum spectroscopy, both the scope of accessible methods and samples must be expanded. HED’s current main limitation, shared with all other entangled-pair spectroscopy, is the low photon flux and coincidence rates from entangled-photon sources. HED is therefore currently not competitive for routine lifetime measurements. However, it can provide a foundation for extending quantum-light spectroscopy beyond simple intensity correlations, opening the possibility to resolve additional energy- and polarization-correlations. Such capabilities would enable more sophisticated approaches for probing quantum systems, with the broader aim of accessing entanglement signatures and unraveling excitonic dynamics in quantum materials [saglamyurek2011broadband]. In light of these requirements, the expansion of entangled-pair spectroscopy from solution-phase organic molecules to solid-state materials is a logical next step. First, the entangled light-matter interaction and emission can be photonically enhanced to increase both the overall brightness and excitation rate per emitter. The latter is needed, for example, to validate theoretical predictions of inter-pair delay-dependent two-photon excitation i.e., electronic pathway selectivity [fujihashi2024pathway].
@@
Here, we demonstrate HED in solid-state materials at low temperatures, specifically thin-films of near-infrared (NIR) emitting colloidal InAs/ZnSe quantum dots (QDs). QDs offer a rich platform for exploring entangled-pair light–matter interactions, thanks to their tunable electronic fine structure, phonon coupling, and electron–hole radiative and non-radiative rates [bera2010quantum]. To our knowledge, our results represent the first demonstration of entangled-photon heralded emission detection in a solid-state material, distinct from prior two-photon excitation approaches using entangled photons [pandya2024towards]. QD’s processability and proven integration with nanophotonic structures may help overcome brightness limitations through enhancement of the absorption and emission [chen2021integration], and serve as a model system for entangled-photon non-linear spectroscopy.
@@ 1 Result
We developed an improved HED apparatus capable of operating over variable solid-state sample temperatures, as shown in Fig. 1a. A detailed description of the setup components is provided in the Supporting Information. Briefly, entangled photon pairs are generated via SPDC by pumping a periodically poled potassium titanyl phosphate (ppKTP) crystal with a 405 nm CW laser under the type-II phase-matching condition. The generated pairs exhibit a bandwidth of $\sim$2 nm with spectral overlap consistent with degenerate SPDC (Fig. 1b). A Hong–Ou–Mandel (HOM) visibility of 86% confirms the high indistinguishability of the photon pairs, in agreement with previous reports (Fig. S1a) [osorio2013purity]. The two photons are separated by a fiber-coupled polarization beam splitter (PBS), yielding a heralding arm and a sample arm. The heralding photons are directly detected by an SNSPD, while the sample-arm photons are sent through a 50:50 beamsplitter into a cryostat with a 0.9 numerical aperature (NA) cryo-objective to excite the sample. The resulting emission is collected by the same objective, filtered from the excitation photons, and coupled via a fiber collimator into a second SNSPD. The use of SNSPDs is a notable feature of our apparatus due to their high photo-detection efficiency in the NIR ($\sim$80%), minimal dark counts (<<10 Hz), and low dead time ($\leq$35 ns). Their additionally low timing jitter (31 ps and 50 ps) provides an overall narrow ($\sim$72 ps) instrument response function (IRF) for coincidence counts without sample (Fig. 1c). The IRF width is mainly determined by the convolution of the timing jitter of the SPDC source, SNSPDs, and the single-photon counting module (19 ps).
Colloidal InAs/ZnSe core–shell QDs were selected as the model system for demonstrating HED [li2025shell]. These nanocrystals are promising NIR emitters [jalali2022indium] and have been widely incorporated into photonic devices such as photoconductors [zhao2019general], phototransistors [kim2023ultrasensitive], and photodiodes [choi2021ligand, sun2022fast]. A key advantage of colloidal InAs QDs is their compatibility with planar photonic architectures: they can be directly deposited onto a variety of substrates and integrated with on-chip photonic structures [elshaari2020hybrid], offering a practical route toward scalable quantum photonic technologies relevant to optical communication and quantum information processing.
@@
Our synthesized type-I heterostructured InAs/ZnSe core–shell nanocrystals display high (approximately 59%) ensemble photoluminescence quantum yield (PLQY) in solution, consistent with previously reported values for this material system [li2025shell]. This PLQY is substantially higher than of organic dyes in the NIR [rurack2011fluorescence], rendering NIR-QDs a natural choice for HED with high signal strengths. Room temperature QDs ensemble absorption and emission spectra reveal a low prominence excitonic shoulder around 905 nm based on Elliott model fitting (Fig. S2g) and a redshifted, broad emission around 984 nm, shown in Fig. 1d. We attribute the absence of narrower features to inhomogeneous broadening due to size- and shape-inhomogeneity, which is common even for the currently best InAs QDs [jalali2022indium, gazis2023colloidal]. The transmission electron microscopy (TEM) image, as shown in the inset of Fig. 1d, confirms the tetrahedral QD shape [zhu2023boosting] and a mean particle size of 8.2 $\pm$ 2.2 nm, with a core size of 4.3 nm and shell thickness of 2.0 nm (Fig. S2f). This size is only a fraction of the InAs bulk Bohr radius of around 35 nm [kan2003synthesis], rendering our QDs well within the strong confinement regime, where size dispersion has a high impact on spectral broadening [jalali2022indium].
For HED measurements, we prepared emissive QD thin films by drop-casting onto gold-coated coverslips (Fig. 1a). Compared with conventional spin-coating, drop-casting produces thicker, higher-optical-density films that result in stronger photoluminescence, thereby improving signal (see Supporting Information).
@@
Figure 2: Data analyses in heralded emission detection. (a) In heralded emission detection (HED), photon pairs are split into a heralding arm (blue) and a sample arm (orange). Pairs can be classified as coincidence pairs, which are time-correlated and produce bunching in the $g^{(2)}(\tau)$, or non-coincidence pairs, which contribute uncorrelated background. The sample emission therefore possesses a similar statistics with emission photon from both twins (red) and accidental pairs (light red). (b) Intensity traces of the heralding arm and the emission signal during a 20h measurement at 4.7 K, demonstrating stable down-conversion over a ten-hour period. (c) Normalized cross-correlation, $g^{(2)}(\tau)$, between the heralding arm and the emission at 4.7 K. The orange-shaded region corresponds to contributions from correlated heralding–emitted photon pairs (entangled two-photon down-conversion), while the purple region represents uncorrelated background with $g^{(2)}(\tau)=1$. (d) Schematic for the $g^{(2)}$ correlation analysis of time-tagged HED data. Delaying one sequence relative to the other only shifts the bunching peak without changing its shape. (e) Schematic of the effect of adopting a ’start–stop’ alternative from time-correlated single-photon counting (TCSPC) to HED, introducing distortions.
@@
Fig. 2 illustrates the concept, stability of emission, and different approaches for analyzing HED data. Three photon “arms” are considered in the measurement: the heralding arm, the sample arm used to excite the sample, and the emission arm, which collects photons emitted by the sample following excitation. SPDC sources are not ideal [ramelow2013highly]; in addition to generating true coincidence photon pairs, non-coincidence pairs are also produced, as shown in Fig. 2a. Consequently, the emitted photons originate from both non-coincidence and coincidence pairs. The former produces accidental heralding-emission coincidences. The sample emission count rates remained stable over the course of the experiment ($\sim$15–18 hours), confirming the sample’s stability under low excitation fluences (see Fig. 2b for 4.7 K). At 4.7 K, the heralding arm consistently recorded an average count rate of $\sim 3.2\times 10^{6}$ counts per second (cps). Assuming the sample arm exhibits a similar rate, and accounting for losses through optical elements, the number of excitation photons reaching the sample is estimated to be $1.12\times 10^{6}$, corresponding to a photon power of $\sim$275 fW. Emitted photons were collected at a rate of $\sim$89 cps (see Supporting Information for details).
@@
We now analyze correlations (see Supporting Information for details) between heralding and emission photons, extending approaches previously applied to molecular dyes at room temperature [harper2023entangled, li2023single, eshun2023fluorescence, gabler2025benchmarking, alvarez2025correlated]. Leveraging the short 72 ps IRF of our HED setup, we resolve the normalized cross-correlation, $g^{(2)}(\tau)$, between emission and heralding photons (Fig. 2c). As expected, a distinct bunching peak >1>1 at $\tau=0$ indicates preservation of the bunched statistics of twin pairs in the heralded emission. The background with $g^{(2)}(\tau)=1$ results from accidental coincidence events. The decay of $g^{(2)}(\tau)>1$ with $\tau$ measures the effective delay between heralded photon absorption and emission, i.e. the spontaneous emission lifetime [harper2023entangled]. We confirm the near-equivalency of HED lifetimes of our QDs and TCSPC, as shown in the Supporting Information. Small deviations are most likely due to differences in the excitation wavelength or small deviations in setup alignments between different excitation sources. Interestingly, we identify a consistent dip of $g^{(2)}<1$ for $\tau<0$ (Fig. 2c, inset). Such anti-correlations would be consistent with time reversal symmetry breaking, but likely have a more trivial origin. We assign the origin of this dip to a non-obvious artifact from detector deadtime. Coincidences with $\tau<0$ are from emitted photons leading heralding photons. A leading emitted photon has a larger than random probability of being led by a second, even earlier heralding photon. The detection of this second, earliest, heralding photon turns the heralding arm SNSPD dark for the deadtime (<<20 ns). The number of following emission/heralding coincidences within the deadtime is thus reduced, producing the dip in $g^{(2)}(\tau<0)$. This artifact may be misinterpreted or may go unnoticed, especially for longer deadtime detectors. While the above correlation analysis is established, we like to point out that translation of the “start-stop” paradigm used in pulsed-laser TCSPC would require correction in post-processing and is more prone to distortion from pathlength differences. We clarify this in Fig. 2d-e.
@@
The start–stop algorithm selects a photon on one channel as “start” and a photon on the other channel as “stop,” storing their time differences in a histogram. By contrast, the correlation algorithm accumulates all time differences between detection events on the two channels into a histogram, treating every click as both “start” and “stop,” and calculating both positive and negative delays. Using a start–stop scheme in HED can introduce distortions, since the overall number of pairs along the lifetime not only depends on the heralding ratio and entanglement time, but also the waiting time distribution of the accidental pairs. This can produce spurious accidental coincidences that are difficult to separate from the lifetime decay. Timing misalignment between emitted photon and their heralding photon can further deform the decay dynamics (Fig. 2d). The $g^{(2)}$ correlation approach avoids these complications, as it does not impose fixed start/stop assignments on the photons. Any time delay between the heralding and emission arms merely shifts the correlation peak without distorting its decay dynamics (Fig. 2e). Waiting time distributions between accidental counts are naturally factored out in the $g^{(2)}$.
@@
One shortcoming of many SNSPDs is their single-mode fiber access [kim2011efficient], causing substantial insertion losses. If high count rates rather than minimal timing jitter are desirable in HED, large area APDs are preferred. Fig. 3a–c showcases HED results acquired across different temperatures using both an SNSPD (Fig. 3a) and a large-area APD (Fig. 3b-c). Although the employed APD exhibited high dark counts ($\sim$1848 cps), they contribute only to accidental coincidences and do not affect the shape of the bunching peak (see Supporting Information for details). We consistently observed total sample emission count rates of $\sim$2000 cps when subtracting accidental dark counts. Sample 1, with a quantum dot size of 8.2 nm, and Sample 2, with a size of 6.0 nm, showed similar estimated brightness despite their differing lifetimes. These detected count rates led to an increased external quantum efficiency (EQE) for the film i.e., the ratio of detected emission photons to heralded sample arm photons (Fig. 3f).